An ecoli bacteria divides every 20 minutes. After 120 minutes how many ecoli bacteria are there? Solve algebraically.
Answer by stanbon(75887)

An ecoli bacteria divides every 20 minutes. After 120 minutes how many ecoli bacteria are there? Solve algebraically.
------------------------
A(t) = Ao (2)^(t/20)
Assuming you start with 1 bacteria you get:
A(120)=1(2)^(120/20)
=2^6=64
That is the number generated in the last 20 minutes.
---------------
The total number = 1+2+4+....+64
---------------
This is a geometric series with seven terms, a(1)=1, r=2:
S(7)= [r^(8)-1}/[r-1]= 127
--------------
Cheers,
Stan H.
-----------